OpenStudy (anonymous):

whats up guys , can someone explain me how to solve this type of fractions (1+1/x) / ( 1-1/x^2) ,

OpenStudy (anonymous):

i know that the answer is x/x-1 ,but i dont know how to do it

hero (hero):

\[\dfrac{1 + \dfrac{1}{x}}{1 - \dfrac{1}{x^2}}\]

hero (hero):

Is that what you have to begin with?

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

thats the fraction

OpenStudy (anonymous):

do u know how to solve it ,or the name of this fractions , cause i dont know how to do it and i want to know , thank u for answer my question

hero (hero):

\[\dfrac{\dfrac{x}{x} + \dfrac{1}{x}}{\dfrac{x^2}{x^2} - \dfrac{1}{x^2}}\] \[\dfrac{\dfrac{x + 1}{x}}{\dfrac{x^2 - 1}{x^2}}\] \[\dfrac{x + 1}{x} \div \frac{x^2 - 1}{x^2}\] \[\dfrac{x + 1}{x} \times \frac{x^2}{x^2 - 1}\] \[\frac{x + 1}{x} \frac{x^2}{(x + 1)(x - 1)}\] \[\frac{\cancel{x + 1}}{\cancel{x}} \frac{x^\cancel{2}}{\cancel{(x + 1)}(x - 1)}\] \[\frac{x}{x - 1}\]

OpenStudy (anonymous):

thank u , do u know the name of this fractions

hero (hero):

What do you mean? I think maybe complex fractions.

OpenStudy (anonymous):

na dont worry and thank u i got it